# Quadratic exchange equations for Coxeter matroids

### v1

@@ A.2 Explicit equations: Type $B_{n}$
It was shown in the proof of Lemma 4.13 that due to the mod $4$ conditions, $2|M|$ cannot be m2m-2 or $m+2$. Thus it makes no difference to the theorem whether those are listed when defining ${\mathcal{M}}_{N,L}$. However, explicitly removing $\{m-2,m-1,m,m+1,m+2\}$ from ${\mathcal{M}}_{N,L}$ clearly demonstrates that that there are no equations supported on antipodes of 0-, 11- and $2$-cubes.
@@
To prove Theorem 4.9, it is sufficient to prove that the quadrics $\{p^{M}_{N,L}\}$ span the summand $S^{2}(\mathcal{S}^{\vee})/V_{2\lambda}$. Note that the Casimir $\Omega$ has different eigenvalues on each irreducible summand of $S^{2}(\mathcal{S}^{\vee})$. The lemma below calculates each of these eigenvalues, producing a basis for $S^{2}(\mathcal{S}^{\vee})$ which splits into bases for each irreducible summand. We then finish the proof of Theorem 4.9 by deducing which subset of this basis is a basis for $S^{2}(\mathcal{S}^{\vee})/V_{2\lambda}$.
For easier notation, recall as in Definition 4.10 that
$K,\overline{K}]=|\overline{K}^{c}|\left(|N|+|L|\right)+\ell\left({\overline{K}^{c}}^{T}\circ\overline{K}\right)+n|\overline{K}|+\ell\left(L\circ\left({K}\!\setminus\!N\right)\right)+\ell\left({N}^{T}\circ{K}\!\setminus\!N\right)$
where $(K,\overline{K})$ is an ordered antipode in $Q_{N,L}$.
@@
Proof of Theorem 4.8.
Lemma A.11 gives a basis of $\Omega$-eigenvectors for $S^{2}(\mathcal{S}^{\vee})$, and this basis splits into a basis for each irreducible submodule. To prove Theorem 4.8, it only remains to remove the basis elements that span $V_{2\lambda}$. Note that $2\lambda=(1,\ldots,1)$ which is the highest weight for ${\textstyle{\bigwedge}}^{n}V\cong{\textstyle{\bigwedge}}^{n+1}V$. One can see this by noticing that all of the weights of $V$ are $(0,\dots,\pm 1,\dots,0)$ or $(0,\ldots,0)$ with multiplicity one. The weights of ${\textstyle{\bigwedge}}^{k}V$ are all sums of $k$ different weights of $V$. Notice, one must use either $n$ or $n+1$ weights from $V$ to sum to $2\lambda$. Therefore, we must remove the basis elements given in Lemma A.11 that originate from ${\textstyle{\bigwedge}}^{n}V$ or ${\textstyle{\bigwedge}}^{n+1}V$, only one of which can occur for each n. The $z$ and $z^{\prime}$ that originate in ${\textstyle{\bigwedge}}^{n}V$ are such that $|N|+2|M|+|L|=n$ and $|N^{\prime}|+2|M^{\prime}|+|L^{\prime}|=n-1$. Likewise, $z$ and $z^{\prime}$ that originate in ${\textstyle{\bigwedge}}^{n+1}V$ are such that $|N|+2|M|+|L|=n+1$ and $|N^{\prime}|+2|M^{\prime}|+|L^{\prime}|=n$. Our basis of $S^{2}(\mathcal{S})$ constructed in Lemma A.11 are the $p^{M}_{N,L}$ such that
@@
To obtain a basis for the subspace of $S^{2}(\mathcal{S}^{\vee})$ orthogonal to $V_{2\lambda}$, we remove the $p^{M}_{N,L}$ such that $|N|+2|M|+|L|\in\{n-1,n,n+1\}$. By the mod $4$ conditions, $|N|+2|M|+|L|$ is never equal to n2n-2 or $n+2$: this is shown in the proof of Lemma 4.13. Thus we find a basis for the Lichtenstein embedding equations is the set of $p^{M}_{N,L}$ such that
@@ A.3 Equations from Lichtenstein: Type $D_{n}$
We now proceed to find a basis for the Lichtenstein embedding equations for type $D_{n}$ with its minuscule representation $\mathcal{S}_{+}$. The organisation and procedure is identical to type $B_{n}$ above. We focus on $\mathcal{S}_{+}$ since the two varieties associated to $\mathcal{S}_{+}$ and $\mathcal{S}_{-}$ are the two connected components of the isotropic Grassmannian and are diffeomorphic as varieties, and furthermore the weights of $\mathcal{S}^{+}$ and $\mathcal{S}^{-}$ are isomorphic as (transitive) sets with a $W$-action. So for our purposes these cases are equivalent. Compare the description for the Type $D_{n}$ Lichtenstein embedding equations in [chirivi2013pfaffians, Corollary 16], whose presentation uses shuffle relations.
@@
Here we implicitly assume $I$ is ordered $i_{1}<\cdots<i_{2k}$ for this basis and $|I|$ is even. Similarly we model the module $\mathcal{S}_{-}$ with $\bigwedge^{\mathrm{odd}}F$. Note that $v_{I}$ is a basis for the weight space with +1/2+1/2 in the $i^{\rm th}$ position if $i\in I$ and 1/2-1/2 otherwise. Also, $\mathcal{S}_{+}$ has weights with an even number of positive and negative entries and $\mathcal{S}_{-}$ has an odd number of positive and negative entries. Let $\mathcal{S}=\mathcal{S}_{+}\oplus\mathcal{S}_{-}$. Then the $\mathfrak{so}_{2n}$-representation is precisely the restriction of the $\mathfrak{so}({2n+1})$-representation $\mathcal{S}$ considered in Section A.1. Therefore the isomorphism of $\mathfrak{so}({2n+1})$-representations considered induces an isomorphism of $\mathfrak{so}({2n})$-representations. Since $\mathcal{S}=\mathcal{S}_{+}\oplus\mathcal{S}_{-}$ we have
@@
Theorem 4.8 shows that $S^{2}(\mathcal{S}^{\vee})$ is spanned by equations $p^{M}_{N,L}$. Thus to find the Lichtenstein embedding equations, we need to find the subspace of the span of these equations that spans $S^{2}(\mathcal{S}_{+}^{\vee})$.
See 4.18
@@
Note that $\mathcal{S}$ has a basis given by ${\mathcal{Q}}_{n}$ and a basis for $\mathcal{S}_{\pm}$ is given by the even resp. the odd demicube ${\mathcal{D}}_{n}^{\pm}$. The space $S^{2}(\mathcal{S}^{\vee})$ can be considered as the space of symmetric $2$-forms on $\mathcal{S}$. The space $S^{2}(\mathcal{S}_{+}^{\vee})$ is the space of symmetric $2$-forms on $\mathcal{S}_{+}$ and can be described as all symmetric $2$-forms on $\mathcal{S}$ that vanish on $\mathcal{S}_{-}$, equivalently vanishing on ${\mathcal{D}}_{n,-}$. Thus the space of Lichtenstein embedding equations of ${\mathbb{G}}/{\mathbb{P}}_{\alpha_{n}}$ for type $D_{n}$ is precisely the subspace of Lichtenstein embedding equations of type $B_{n}$ that are supported only on the even demicube and not the odd demicube. An identical argument works with ${\mathbb{G}}/{\mathbb{P}}_{\alpha_{n-1}}$ and $\mathcal{S}_{-}$. ∎

### v2

@@ A.2.1 A first basis of equations
It will be shown in the proof of Lemma A.16 that due to the mod $4$ conditions, $2|M|$ cannot be m2m-2 or $m+2$. Thus it makes no difference to the theorem whether those are listed when defining ${\mathcal{M}}_{N,L}$. However, explicitly removing $\{m-2,m-1,m,m+1,m+2\}$ from ${\mathcal{M}}_{N,L}$ clearly demonstrates that that there are no equations supported on antipodes of 0-, 11- and $2$-cubes.
@@
To prove Theorem A.10, it is sufficient to prove that the set of quadrics $\{p^{M}_{N,L}\}$ span the summand $S^{2}(\mathcal{S}^{\vee})/V_{2\lambda}$. Note that the Casimir $\Omega$ has different eigenvalues on each irreducible summand of $S^{2}(\mathcal{S}^{\vee})$. The lemma below calculates each of these eigenvalues, producing a basis for $S^{2}(\mathcal{S}^{\vee})$ which splits into bases for each irreducible summand. We then finish the proof of Theorem A.10 by deducing which subset of this basis is a basis for $S^{2}(\mathcal{S}^{\vee})/V_{2\lambda}$.
Definition A.12.
Given an ordered antipode $(K,\overline{K})$, we define
Recall that $N$ and $L$ are determined by $(K,\overline{K})$.
@@
Proof of Theorem A.10.
Lemma A.13 gives a basis of $\Omega$-eigenvectors for $S^{2}(\mathcal{S}^{\vee})$, and this basis splits into a basis for each irreducible submodule. To prove Theorem A.10, it only remains to remove the basis elements that span $V_{2\lambda}$. Note that $2\lambda=(1,\ldots,1)$ which is the highest weight for ${\textstyle{\bigwedge}}^{n}V\cong{\textstyle{\bigwedge}}^{n+1}V$. One can see this by noticing that all of the weights of $V$ are $(0,\dots,\pm 1,\dots,0)$ or $(0,\ldots,0)$ with multiplicity one. The weights of ${\textstyle{\bigwedge}}^{k}V$ are all sums of $k$ different weights of $V$. Notice, one must use either $n$ or $n+1$ weights from $V$ to sum to $2\lambda$. Therefore, we must remove the basis elements given in Lemma A.13 that originate from ${\textstyle{\bigwedge}}^{n}V$ or ${\textstyle{\bigwedge}}^{n+1}V$, only one of which can occur for each n. The $z$ and $z^{\prime}$ that originate in ${\textstyle{\bigwedge}}^{n}V$ are such that $|N|+2|M|+|L|=n$ and $|N^{\prime}|+2|M^{\prime}|+|L^{\prime}|=n-1$. Likewise, $z$ and $z^{\prime}$ that originate in ${\textstyle{\bigwedge}}^{n+1}V$ are such that $|N|+2|M|+|L|=n+1$ and $|N^{\prime}|+2|M^{\prime}|+|L^{\prime}|=n$. Our basis of $S^{2}(\mathcal{S})$ constructed in Lemma A.13 are the $p^{M}_{N,L}$ such that
@@
To obtain a basis for the subspace of $S^{2}(\mathcal{S}^{\vee})$ orthogonal to $V_{2\lambda}$, we remove the $p^{M}_{N,L}$ such that $|N|+2|M|+|L|\in\{n-1,n,n+1\}$. By the mod $4$ conditions, $|N|+2|M|+|L|$ is never equal to n2n-2 or $n+2$: this is shown in the proof of Lemma A.16. Thus we find a basis for the Lichtenstein embedding equations is the set of $p^{M}_{N,L}$ such that
@@
A.2.2 Minimising the support
We now produce the type $B$ quadratic embedding equations $F_{I,J}^{(B)}$ of Theorem 4.8. To work with these equations it will be convenient to rephrase and reindex them. Recall the notation for subcubes $Q_{N,L}$ and ordered antipodes $(K,\overline{K})$ from Section A.2.1. For such an ordered antipode $(K,\overline{K})$ with $\overline{K}=K\mathbin{\triangle}S$
In Lemma A.13 we presented the Lichtenstein embedding equations as
for all $M$ with $2|M|\in{\mathcal{M}}_{N,L}$. The proof of Theorem A.10 shows that $p^{M}_{N,L}$ is symmetric under exchanging $K$ and $\overline{K}$, but in general $K,\overline{K}]$ differs from $\overline{K},K]$, and the two can even have different parities.
To prove Theorem 4.8 it suffices to show that the span of the quadratic embedding equations $\{\mathrm{ex}_{(K,K\mathbin{\triangle}S)}\}$ is equal to the span of the Lichtenstein embedding equations $\{p^{M}_{N,L}\}$. The rest of this subsection is dedicated to proving the equality of these subspaces of quadrics. Corollary A.18 is the forward inclusion and Lemma A.19 is the reverse.
To compare the quadratic embedding equations with the Lichtenstein embedding equations, we must rewrite them with coefficients in terms of $K,\overline{K}]$. This is the content of Corollary A.15. We first give a technical lemma.
Lemma A.14.
Proof.
We use Definition A.12 to calculate the difference $K,\overline{K}]-[K\mathbin{\triangle}i,\overline{K}\mathbin{\triangle}i]\mod 2$ term by term. We first observe that
Repeating this argument for the other permutations in Definition A.12, we get
Corollary A.15.
Up to a global sign, we have
Proof.
If we multiply $\mathrm{ex}_{(K,K\mathbin{\triangle}S)}$ from Definition 4.7 by the sign $(-1)^{n+[K,K\mathbin{\triangle}S]}$, the claim immediately follows from Lemma A.14. ∎
We next introduce the variables $y_{(K,\overline{K})}=(-1)^{[K,\overline{K}]}x_{K}x_{\overline{K}}$. We emphasise that as $(-1)^{[K,\overline{K}]}$ may differ from $(-1)^{[\overline{K},K]}$, the variables $y_{(K,\overline{K})}$ and $y_{(\overline{K},K)}$ may differ. By Corollary A.15, in these variables, the quadratic embedding equations become
and the Lichtenstein embedding equations become
This change of variables helps us prove the following.
Lemma A.16.
Proof.
We next define our intermediary polynomials $\chi_{b}$, as follows:
$\chi_{b}=\sum_{|M|=b}p^{M}_{\varnothing,S^{c}}(y)\,.$
We prove (62) by instead considering a more general function
$S(m,a,c)\coloneqq\sum_{b\equiv c\bmod 2}\frac{1}{2}(c-b)\chi_{b}(m,a)$
where $c\in{\mathbb{Z}}$ and
$\chi_{b}(m,a):=\sum_{k}(-1)^{k}\binom{a}{k}\binom{m-a}{b-k}\,$.
$S(m,0,c)=\frac{1}{4}(S^{+}(m,c)+(-1)^{c}S^{-}(m,c)).$
When $a>0$ we have
$=\chi_{b}(m-1,a-1)-\chi_{b-1}(m-1,a-1)$
We finish by doing this check. When $c$ is even,
$S(3,a,c)$ $=\frac{c}{2}\chi_{0}(3,a)+\frac{c-2}{2}\chi_{2}(3,a)$
$=\frac{c}{2}+\frac{c-2}{2}(2a^{2}-6a+3).$
The evaluations of $a$ appearing in (63) are
$S(3,0,c)$ $=2c-3=2^{3-2}(c-\textstyle\frac{3}{2}),$
$S(3,1,c)$ $=1=2^{3-3}.$
In the same way, when $c$ is odd,
$S(3,a,c)$ $=\frac{c-1}{2}\chi_{0}(3,a)+\frac{c-3}{2}\chi_{2}(3,a)$
$=\frac{c-1}{2}(-2a+3)+\frac{c-3}{2}(-\frac{4}{3}a^{3}+6a^{2}-\frac{20}{3}a+1),$
and the evaluations in (63) are again
$S(3,0,c)$ $=2c-3,$
$S(3,1,c)$ $=1$
as desired. ∎
Having shown that every $\operatorname{ex}_{(\varnothing,S)}$ is in the span of the Lichtenstein embedding equations, the next step is to generalise to all quadratic embedding equations. Every antipode of every subcube of the $n$-cube is in the orbit of an antipode $\{\varnothing,S\}$ under reflections in the coordinate hyperplanes, so our strategy is essentially to apply these reflections. But we cannot literally act by symmetries of the cube; the symmetries we have are those of the spin representation $\mathrm{Spin}(2n+1)$. To introduce these, we resume the notation relating to the vector representation $V={\mathbb{C}}^{2n+1}$ of $\mathfrak{so}_{2n+1}$ from Appendix A.1.
$\frac{1}{\sqrt{2}}(x_{i}+v_{i})w\in\mathrm{Spin}(2n+1)\,.$
Thus in both cases, in the notation of Section 2.2, we have
Since $g_{i}$ is a group element, we have
$g_{i}(x_{A}x_{B})=g_{i}(x_{A})g_{i}(x_{B})\,.$
Lemma A.17.
Proof.
For any three sets $A,B,C$, we have the associativity relation
Thus, for any $A\subseteq[n]$, we have
Corollary A.18.
Proof.
This follows from Lemma A.17 because the span of this set of $p^{M}_{N,L}$ is closed under the action of the $g_{h}$. ∎
Lemma A.19.
Proof.
$p^{M}_{N,L}[x_{I}x_{\overline{I}}]=(-1)^{|I^{c}\cap M|+[I,\overline{I}]}.$
Note that
Equivalently,
Hence, we find
We also recall from Lemma A.14 that
$=(-1)^{|I^{c}\cap M|+[I,\overline{I}]}\,.$
Proof of Theorem 4.8.
Corollary A.18 and Lemma A.19 prove that the span of the quadratic embedding equations $\{\mathrm{ex}_{(K,K\mathbin{\triangle}S)}\}$ is equal to the span of the Lichtenstein embedding equations $\{p^{M}_{N,L}\}$. As $\{p^{M}_{N,L}\}$ cut out ${\mathbb{G}}/{\mathbb{P}}\subset\operatorname{proj}(\mathcal{S})$ by Theorem A.10, our theorem is proved. ∎
@@ A.3 Equations from Lichtenstein: Type $D_{n}$
We now proceed to find a basis for the Lichtenstein embedding equations for type $D_{n}$ with its minuscule representation $\mathcal{S}_{+}$. The organisation and procedure is identical to type $B_{n}$ above. We focus on $\mathcal{S}_{+}$ since the two varieties associated to $\mathcal{S}_{+}$ and $\mathcal{S}_{-}$ are the two connected components of the isotropic Grassmannian and are diffeomorphic as varieties, and furthermore the weights of $\mathcal{S}_{+}$ and $\mathcal{S}_{-}$ are isomorphic as (transitive) sets with a $W$-action. So for our purposes these cases are equivalent.
@@
Here we implicitly assume $I$ is ordered $i_{1}<\cdots<i_{2k}$ for this basis and $|I|$ is even. Similarly we model the module $\mathcal{S}_{-}$ with $\bigwedge^{\mathrm{odd}}F$. Note that $v_{I}$ is a basis for the weight space with +1/2+1/2 in the $i^{\rm th}$ position if $i\in I$ and 1/2-1/2 otherwise. Also, $\mathcal{S}_{+}$ has weights with an even number of positive and negative entries and $\mathcal{S}_{-}$ has an odd number of positive and negative entries. Let $\mathcal{S}=\mathcal{S}_{+}\oplus\mathcal{S}_{-}$. Then the $\mathfrak{so}_{2n}$-representation is precisely the restriction of the $\mathfrak{so}_{2n+1}$-representation $\mathcal{S}$ considered in Section A.1. Therefore the isomorphism of $\mathfrak{so}_{2n+1}$-representations considered induces an isomorphism of $\mathfrak{so}_{2n}$-representations. Since $\mathcal{S}=\mathcal{S}_{+}\oplus\mathcal{S}_{-}$ we have
@@
Theorem A.10 shows that $S^{2}(\mathcal{S}^{\vee})$ is spanned by equations $p^{M}_{N,L}$. Thus to find the Lichtenstein embedding equations, we need to find the subspace of the span of these equations that spans $S^{2}(\mathcal{S}_{+}^{\vee})$.
See 4.10
@@
Note that $\mathcal{S}$ has a basis given by ${\mathcal{Q}}_{n}$ and a basis for $\mathcal{S}_{+}$ is given by the even demicube ${\mathcal{D}}_{n}^{+}$. The space $S^{2}(\mathcal{S}^{\vee})$ can be considered as the space of symmetric $2$-forms on $\mathcal{S}$. The space $S^{2}(\mathcal{S}_{+}^{\vee})$ is the space of symmetric $2$-forms on $\mathcal{S}_{+}$ and can be described as all symmetric $2$-forms on $\mathcal{S}$ that vanish on $\mathcal{S}_{-}$, equivalently vanishing on ${\mathcal{D}}_{n,-}$. Thus the space of Lichtenstein embedding equations of ${\mathbb{G}}/{\mathbb{P}}_{\alpha_{n}}$ for type $D_{n}$ is precisely the subspace of Lichtenstein embedding equations of type $B_{n}$ that are supported only on the even demicube and not the odd demicube. An identical argument works with ${\mathbb{G}}/{\mathbb{P}}_{\alpha_{n-1}}$ and $\mathcal{S}_{-}$. ∎
As with the odd-dimensional case, the Spin module $\mathcal{S}={\textstyle{\bigwedge}}^{\bullet}F$ is a $\operatorname{Cl}(V)$ module via the action
$\operatorname{Cl}(V)\to\operatorname{End}(\mathcal{S}),\quad v\in F\mapsto\sqrt{2}\epsilon_{v}\,,\,x\in F^{*}\mapsto\sqrt{2}\iota_{x}\,.$
Note that $\epsilon_{v}$ and $\iota_{x}$ exchange $\mathcal{S}_{+}={\textstyle{\bigwedge}}^{\mathrm{even}}F$ and $\mathcal{S}_{-}={\textstyle{\bigwedge}}^{\mathrm{odd}}F$, hence $\mathcal{S}_{+}$ and $\mathcal{S}_{-}$ are $\operatorname{Cl}^{\mathrm{even}}(V)$ submodules. To end, we utilise this Clifford structure to demonstrate that the spin representation of $\operatorname{\mathrm{Spin}}(2n+1)$ is isomorphic to either of the half spin representations of $\operatorname{\mathrm{Spin}}(2n+2)$. As discussed in Remark 4.12, this gives us an isomorphism between the type $B_{n}$ embedding equations and the type $D_{n+1}$ embedding equations.
Let $V_{2n+2}$ be of dimension $2n+2$ with maximal isotropic spaces $F_{n+1}$ and $F^{*}_{n+1}$ such that we have $V_{2n+2}=F_{n+1}\oplus F^{*}_{n+1}$. As described earlier, the spin module $\mathcal{S}_{2n+2}={\textstyle{\bigwedge}}^{\bullet}F_{n+1}$ splits into the two irreducible half-spin representations of $\operatorname{\mathrm{Spin}}(2n+2)$, namely $\mathcal{S}_{2n+2,+}={\textstyle{\bigwedge}}^{\mathrm{even}}F_{n+1}$ and $\mathcal{S}_{2n+2,-}={\textstyle{\bigwedge}}^{\mathrm{odd}}F_{n+1}$, and these are both $\operatorname{Cl}^{\mathrm{even}}(V_{2n+2})$ modules. Let $v_{1},\ldots,v_{n+1}$ be a basis for $F_{n+1}$, with dual basis $x_{1},\ldots,x_{n+1}$ in $F^{*}_{n+1}$. Let $F_{n}$ and $F^{*}_{n}$ be spaces spanned by $v_{1},\ldots,v_{n}$ and $x_{1},\ldots,x_{n}$ respectively and define $V_{2n+1}=F_{n}\oplus F^{*}_{n}\oplus\textrm{Span}\{w\}$ where $w=\sqrt{\frac{-1}{2}}(v_{n+1}-x_{n+1})$. Recall that the spin representation $\mathcal{S}_{2n+1}$ of $\operatorname{\mathrm{Spin}}(2n+1)$ can be modelled as $\mathcal{S}_{2n+1}={\textstyle{\bigwedge}}^{\bullet}F_{n}$. Moreover, it is a $\operatorname{Cl}(V_{2n+1})$ module where $v_{i}$ acts by $\epsilon_{v_{i}}$, $x_{i}$ acts by $\iota_{x_{i}}$ and $w$ acts by $w\cdot v_{I}=(-1)^{|I|}v_{I}$.
Set $e_{2n+2}$ to be $\frac{1}{\sqrt{2}}(x_{n+1}+v_{n+1})$. There is an algebra isomorphism [Meinrenken:2013, pg 79 (3.17)]
$\operatorname{Cl}(V_{2n+1})\cong\operatorname{Cl}^{\mathrm{even}}(V_{2n+2}),\quad v\mapsto\sqrt{-1}ve_{2n+2}\,.$
This allows us to consider $\mathcal{S}_{2n+2,+}$ and $\mathcal{S}_{2n+2,-}$ also as $\operatorname{Cl}(V_{2n+1})$ modules.
Lemma A.20.
There is an isomorphism of $\operatorname{Cl}(V_{2n+1})$ modules
$\phi:{\textstyle{\bigwedge}}^{\bullet}F_{n}\cong{\textstyle{\bigwedge}}^{\mathrm{even}}F_{n+1},\quad v_{I}\mapsto\begin{cases}v_{I}&|I|\text{ even},\\ \sqrt{-1}v_{I\cup\{n+1\}}&|I|\text{ odd}.\end{cases}$
Proof.
We verify the property that for all $v_{I}\in{\textstyle{\bigwedge}}^{\bullet}F_{n}$ and $u$ in the basis $\left\{v_{i},x_{i},w\,\colon\,i\in\{1,\ldots,n\}\right\}$ that
$\sqrt{-1}ue_{2n+2}\cdot\phi(v_{I})=\phi(u\cdot v_{I}),$
where $\cdot$ denotes the Clifford algebra actions on the respective spin modules. Note that $w$ acts on ${\textstyle{\bigwedge}}^{\bullet}F_{n+1}$ by $\sqrt{-1}(\epsilon_{x_{n+1}}-\iota_{v_{n+1}})$ and $e_{2n+2}$ acts by $\epsilon_{v_{n+1}}+\iota_{v_{n+1}}$.
First, suppose $|I|$ is even. Note that we always have $e_{2n+2}\cdot\phi(v_{I})=\epsilon_{v_{n+1}}v_{I}=v_{I\cup\{n+1\}}$, hence
$\sqrt{-1}v_{i}e_{2n+2}\cdot\phi(v_{I})$ $=\sqrt{-1}v_{i}\cdot v_{I\cup\{n+1\}}=\sqrt{-1}v_{i}\wedge v_{I\cup\{n+1\}}=\phi(v_{i}\cdot v_{I})$
$\sqrt{-1}x_{i}e_{2n+2}\cdot\phi(v_{I})$ $=\sqrt{-1}x_{i}\cdot v_{I\cup\{n+1\}}=\phi(x_{i}\cdot v_{I})$
$\sqrt{-1}we_{2n+2}\cdot\phi(v_{I})$ $=\sqrt{-1}w\cdot v_{I\cup\{n+1\}}=\sqrt{-1}(-\sqrt{-1}\iota_{x_{n+1}})v_{I\cup\{n+1\}}=v_{I}=\phi(w\cdot v_{I}).$
Now suppose $|I|$ is odd. Note that we always have $e_{2n+2}\cdot\phi(v_{I})=\sqrt{-1}\iota_{v_{n+1}}v_{I\cup\{n+1\}}=-\sqrt{-1}v_{I}$, hence
$\sqrt{-1}v_{i}e_{2n+2}\cdot\phi(v_{I})$ $=v_{i}\cdot v_{I}=\phi(v_{i}\cdot v_{I})$
$\sqrt{-1}x_{i}e_{2n+2}\cdot\phi(v_{I})$ $=x_{i}\cdot v_{I}=\phi(x_{i}\cdot v_{I})$
$\sqrt{-1}we_{2n+2}\cdot\phi(v_{I})$ $=w\cdot v_{I}=\sqrt{-1}\epsilon_{v_{n+1}}v_{I}=-\sqrt{-1}v_{I\cup\{n+1\}}=\phi(w\cdot v_{I}).$
Thus $\phi$ is a morphism of $\operatorname{Cl}(V_{2n+1})$ modules, and hence $\mathrm{Spin}(2n+1)$ modules. A similar isomorphism exists from ${\textstyle{\bigwedge}}^{\bullet}F_{n}$ to ${\textstyle{\bigwedge}}^{\mathrm{odd}}F_{n+1}$. ∎